OpenStudy (anonymous):

x 1/6 --- = ---- 1/3 x

OpenStudy (***[isuru]***):

hmm.. i think u want to solve this for x... but to do that it would be better to simplify the both sides first .. first consider the left hand side so.. the left hand side simplifies into 3x now if u consider the right hand side  now u will get ur new equation as... \[\frac{ 1 }{ 6x } = 3x \] now do the cross multiplication \[18x^{2} = 1\]\[x^{2} = \frac{ 1 }{ 18 }\]\[x = \pm \sqrt{\frac{ 1 }{ 18 }}\] hope this will help ya :)